in a certain reaction 0.20g of hydrogen gas combines with 3.20g of oxygen gas. if the rmm of the compound is 34.0. find its molecular formula


1
Expert's answer
2021-02-18T07:33:21-0500

n(H) = 0.2 / 1 = 0.2 (mol)

n(O) = 3.20 / 16 = 0.2 (mol)

The ratio H:O = 1:1. If M = 34, then formula is H2O2, hydrogen peroxide.

Answer: H2O2.